# Quantum Down Sampling Filter for Variational Auto-encoder

### v1

@@ Abstract
Variational Autoencoders (VAEs) are essential tools in generative modeling and image reconstruction, with their performance heavily influenced by the encoder-decoder architecture. This study aims to improve the quality of reconstructed images by enhancing their resolution and preserving finer details, particularly when working with low-resolution inputs (16x16 pixels), where traditional VAEs often yield blurred or in-accurate results. To address this, we propose a hybrid model that combines quantum computing techniques in the VAE encoder with convolutional neural networks (CNNs) in the decoder. By upscaling the resolution from 16x16 to 32x32 during the encoding process, our approach evaluates how the model reconstructs images with enhanced resolution while maintaining key features and structures. This method tests the model’s robustness in handling image reconstruction and its ability to preserve essential details despite training on lower-resolution data. We evaluate our proposed down sampling filter for Quantum VAE (Q-VAE) on the MNIST and USPS datasets and compare it with classical VAEs and a variant called Classical Direct Passing VAE (CDP-VAE), which uses windowing pooling filters in the encoding process. Performance is assessed using metrics such as the Fréchet Inception Distance (FID) and Mean Squared Error (MSE), which measure the fidelity of reconstructed images. Our results demonstrate that the Q-VAE consistently outperforms both the Classical VAE and CDP-VAE, achieving significantly lower FID and MSE scores. Additionally, CDP-VAE yields better performance than C-VAE. These findings highlight the potential of quantum-enhanced VAEs to improve image reconstruction quality by enhancing resolution and preserving essential features, offering a promising direction for future applications in computer vision and synthetic data generation.
@@ 1 Introduction
Variational Autoencoders (VAEs) have become foundational tools in generative modeling and image reconstruction, especially in tasks that require efficient and precise feature extraction from high-resolution data [1]. VAEs are part of the broader family of latent-variable models [2], wherein an encoder maps input data to a structured latent space, and a decoder reconstructs the data from this representation. Despite their success in a wide range of applications, VAEs often face challenges when reconstructing high-resolution images, often leading to blurry or low-quality output [3]. These limitations arise primarily from the difficulty of capturing complex high-resolution dependencies, which become especially apparent when dealing with large data sets or complex structures [4]. Recent advancements in quantum computing offer promising avenues for overcoming these challenges by exploiting quantum phenomena such as superposition and entanglement. These properties allow for more efficient data representation and processing, potentially overcoming some of the bottlenecks inherent in classical methods for generative modeling [5]. The integration of quantum circuits into neural networks, particularly within the VAE encoder, has shown potential to improve latent space representations, leading to enhanced feature extraction and more accurate reconstructions [6]. Quantum in VAE represents an example of such hybrid models, which leverage quantum operations to better capture the underlying data structure, making them especially well suited for high-resolution image tasks [5]. Recent advances in generative models have seen the increasing use of low-resolution images, such as 8x8 or 16x16 pixels, as starting points for training high-quality outputs [7]. This approach might seem counterintuitive because of the small size of these images, but it has proven effective for several reasons. Training generative models on low-resolution images allows for faster training times, as smaller images require less computational power and processing, enabling quicker iterations and fine-tuning. In addition, low-resolution images force the model to focus on learning high-level patterns and structures rather than fine-grained details. This simplification helps the model build a strong foundation for understanding the essential features of the data. Once the model has learned to generate meaningful, broad structures at low resolution, it can then be used to upscale or refine the output to a higher resolution, such as 32x32, using techniques like super-resolution or progressive refinement. This method allows the model to generate more detailed and sharper images while avoiding the computational difficulties that come with directly training on higher resolutions like 32x32. Although 32x32-pixel images are commonly used in many applications, including image classification and generative modeling, classical models like VAEs struggle with these resolutions due to limitations in their architecture [8]. VAEs operate by encoding input data into a compressed probabilistic latent space that is then used to reconstruct the input image. However, with 32x32 images, the encoder has to condense a large amount of pixel information into a smaller latent space, often losing fine details in the process. This leads to blurry or artifact-filled reconstructions, as the latent space might not capture the complexity of the image at this resolution. The encoder’s limited capacity to represent detailed features in the latent space becomes particularly problematic for 32x32 images, which are still complex enough to contain important texture, color variation, and structural information. As a result, VAEs can struggle to generate high-quality images from these lower-resolution inputs [9]. The use of 32x32 resolution is prevalent in many machine learning applications, particularly in image classification tasks and generative modeling, as it strikes a balance between computational efficiency and sufficient detail. Datasets like CIFAR-10 and CIFAR-100, which consist of 32x32 pixel images, are standard benchmarks for evaluating generative models and classifiers [10]. Additionally, modern generative models often employ techniques like progressive training, where low-resolution images (8x8 or 16x16) are initially learned, and then progressively refined to higher resolutions, such as 32x32 or beyond. This process stabilizes training and reduces the complexity associated with generating high-resolution outputs [11]. In practice, combining low-resolution training (e.g., 8x8 or 16x16) with upscaling techniques allows generative models to learn more efficiently and handle the complexities of higher-resolution images. By training on smaller images first, the models can focus on understanding the broader patterns in the data, then gradually increase the resolution to produce more detailed outputs. This method has potential to be useful for reducing the computational load and improving the quality of the final images generated. In this work, we introduce a novel Q-VAE that integrates quantum encoding techniques in the encoder while using Convolutional Neural Networks (CNNs) in the decoder. This hybrid architecture leverages the high expressiveness of quantum states to improve feature extraction and mitigate the blurriness often associated with classical VAEs. Unlike traditional quantum generative models, which typically operate on lower-resolution datasets (e.g. 8x8 or 16x16 pixels) [11], our approach is designed to handle higher-resolution inputs, such as 32x32 pixels, allowing the model to produce precise and high-quality image reconstructions for more complex datasets. In the context of quantum-enhanced Variational Autoencoders (Q-VAE), we use datasets MNIST and USPS primarily due to their simplicity and relevance to early-stage exploration of quantum models in image generation. Both are grayscale (black-and-white) datasets, which are less complex than color images, making them ideal for testing new techniques, particularly when working with emerging technologies like quantum computing in machine learning. Since quantum computing, especially its integration into deep learning models, is still in the early stages, focusing on simpler datasets allows for a more controlled environment where we can better understand the impact of quantum techniques on model performance. Furthermore, the datasets used in this study, such as MNIST and USPS (with image resolutions of 28x28 and 16x16 pixels, respectively), are relatively small in size, which is advantageous when exploring the potential of quantum circuits for feature extraction and latent space representation. Their reduced resolution helps lower the computational burden compared to higher-resolution images. To ensure a fair evaluation, the MNIST dataset is first down-sampled from 28x28 to 16x16 pixels, and then, like the USPS dataset, it is up-sampled to 32x32 pixels before being used as input to assess our model’s performance. By focusing on grayscale images, we reduce the complexity of the problem, allowing for clearer insights into how quantum techniques can improve feature extraction and image reconstruction, particularly for simpler tasks like recognizing digits. Once the approach is successfully validated on these datasets, the methods can be extended to more complex and higher-resolution data, such as color images or higher resolutions. Quantitative analysis have been conducted for comparing the effectiveness of the model by using common performance metrics, such as Fréchet Inception Distance (FID) and Mean Squared Error (MSE), which are essential for assessing the quality of the generated images and the latent space representation [17, 18]. Thus, by leveraging these datasets, we can benchmark the Q-VAE against traditional models,to explore that if quantum component brings measurable improvements to the generative process. In summary the following contributions have been made in this paper:
@@
  • New Model: Introduces a quantum-enhanced VAE (Q-VAE), where quantum encoding is applied only in the encoder, simplifying the model and improving feature extraction capabilities.
@@
  • Comparison with Classical Approach: Compares Q-VAE with the classical CDP-VAE, highlighting performance differences due to the inclusion of quantum encoding in the encoder.
@@
  • Simplification: Avoids quantum entanglements or layers in the encoder, focusing solely on an efficient quantum approach for latent space encoding.
@@
  • Reduction in Quantum Trainable Parameters: Minimizes quantum trainable parameters, leading to more efficient model training and better utilization of quantum resources.
@@
  • Potential for Improved Generative Performance: Evaluates whether quantum encoding improves generative quality and latent space representation compared to classical methods.
@@
  • Low to High Resolution Conversion: Investigates whether converting images from low resolution (16x16) to high resolution (32x32) improves model performance.
@@
The structure of this paper is as follows: Section 2 provides a review of related work on classical VAEs and their integration with quantum computing. Section 3 presents the background on Variational Autoencoders (VAEs). In Section 4, the methodology for the proposed quantum-enhanced VAE model is introduced. Section 5 outlines the numerical simulations conducted, while Section 6 discusses the performance analysis. Finally, Section 7 concludes the paper, summarizing the key findings and suggesting potential avenues for future research.
@@ 2 Related Research
In recent years, Variational Autoencoders (VAEs) have garnered significant attention in generative modeling, especially for image reconstruction tasks [19]. Traditional VAEs are praised for their ability to encode data into a structured latent space and regenerate approximations of the input. However, classical VAEs often suffer from blurred reconstructions when dealing with high-resolution images [3], prompting research into improving VAE architectures for better image quality and feature extraction [23, 24, 25]. Several studies have focused on enhancing VAE performance using convolutional architectures [41, 28, 29, 30]. Gatopoulos et al. introduced convolutional VAEs to better capture image features, surpassing the performance of fully connected architectures on high-resolution datasets such as CIFAR-10 and ImageNet32 [26]. Additionally, combining VAEs with adversarial training, such as in Adversarial Autoencoders (AAEs) [31], incorporates discriminator networks to enhance the visual quality of generated images, addressing the blurriness issues inherent in classical VAEs and improving machine learning applications. Moreover, Adversarial VAEs [32, 33] and VAE-GAN hybrids [34] enhance image generation by incorporating adversarial training, leading to sharper images by reducing blurriness [35]. Quantum Machine Learning (QML) has emerged as a promising field, with studies exploring the integration of quantum circuits into generative models. Quantum GANs (QGANs) leverage quantum computing to enhance the data representation and image generation capabilities of GANs. While QGANs are not directly related to VAEs, they represent a quantum-based generative approach that improves data diversity and generation quality [37]. Quantum Variational Autoencoders (Q-VAEs) [38, 39, 40, 41, 42, 43] build on the foundational work by Khoshaman et al. (2018), who proposed a quantum VAE (QVAE) where the latent generative process is modeled as a quantum Boltzmann machine (QBM). Their approach demonstrates that QVAE can be trained end-to-end using a "quantum" lower bound to the variational log-likelihood, achieving state-of-the-art performance on the MNIST dataset [36]. Gircha et al(2021). further advanced the quantum VAE architecture by incorporating quantum circuits in both the encoder, which utilizes multiple rotations, and the decoder [40]. These quantum VAE models primarily focus on leveraging quantum methods in the latent space representation, enhancing the generative capabilities of VAEs. This paper addresses a gap in the literature where quantum techniques are primarily used in hybrid models for the encoder, while the decoder remains classical. Our proposed model is unique in that it leverages quantum encoding with a single rotation gate and measurement solely in the encoder, distinguishing it from other approaches.
@@ 3 Background
Variational Autoencoders (VAEs) are widely recognized as powerful generative models in machine learning. They are based on a probabilistic framework where an encoder network maps input data $\mathbf{x}$ to a latent distribution $q(\mathbf{z}|\mathbf{x})$, and a decoder network reconstructs the data from the latent variables $\mathbf{z}$.
$\mathcal{L}(\theta,\phi)=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\left[\log p(\mathbf{x}|\mathbf{z})\right]-\text{KL}(q(\mathbf{z}|\mathbf{x})\|p(\mathbf{z}))$ (1)
where $\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}$ represents the expectation with respect to the approximate posterior $q(\mathbf{z}|\mathbf{x})$, and $\text{KL}(q(\mathbf{z}|\mathbf{x})\|p(\mathbf{z}))$ is the Kullback-Leibler (KL) divergence between the approximate posterior and the prior distribution $p(\mathbf{z})$ [12]. The reconstruction term encourages the model to generate data points that are similar to the observed data, while the KL divergence regularizes the learned latent variables by pushing the approximate posterior towards the prior distribution. While VAEs have been successful in many domains, including image generation and anomaly detection, they often face significant challenges when applied to high-resolution data. One of the most notable limitations is that traditional encoders, typically implemented as fully connected or convolutional neural networks (CNNs), may fail to capture complex structures in high-resolution datasets, leading to blurred reconstructions. This issue arises from the latent space often not being expressive enough to model the intricate relationships present in high-resolution data, such as large images (e.g., $32\times 32$ pixels or higher) [13].
@@ 3.1 Encoder
A classical VAE encoder is typically implemented as a fully connected network or a Convolutional Neural Network (CNN) within an Artificial Neural Network (ANN), aiming to map input data to a lower-resolution latent space. The primary objective of the encoder is to process the input data $x$, transform it through a series of hidden layers, and output two parameters: the mean $\mu$ and the log variance $\log(\sigma^{2})$, which are used for variational inference [1]. In the case of image data, the input image is first reshaped into a vector and passed through multiple fully connected layers. The transformations performed by these layers help extract important features from the input. For example, given an input image of dimension 1024 (flattened from a $32\times 32$ image), the encoder progressively reduces the resolution’s, often mapping the input image $x_{0}$ into a lower-resolution space $x_{1}$.
3.1.1 ANN as CNN or Fully Connected Layers in Encoders
While ANNs serve as a classic encoder architecture, some models combine Convolution Neural Networks (CNNs) with fully connected layers to enhance the feature extraction process. CNNs are particularly useful for image-based data because they can capture local spatial patterns (e.g., edges, textures) by using convolution layers. These convolution layers progressively down-sample the input image and extract high-level features, while the fully connected layers act to compress these features into a more compact latent representation. In CNN-based encoders, after the convolution layers extract relevant patterns, fully connected layers are used to map the high-resolution feature maps into the latent space. This approach allows the model to better capture spatial dependencies while still benefiting from the representational power of fully connected layers.

### v2

@@ Abstract
Variational Auto-encoders (VAEs) are crucial for generative modeling and image reconstruction, but their performance often degrades with low-resolution input (16x16 pixels), leading to blurred or inaccurate results. This study proposes a hybrid model that combines quantum computing in the VAE encoder (Q-VAE) and convolutional neural networks (CNNs) in the decoder. By upscaling the resolution from 16x16 to 32x32 during encoding, we assess the model’s ability to preserve key features while improving image resolution. We tested the model on the MNIST and USPS datasets, comparing it to VAE refered as classical VAE (C-VAE) and the classical direct-passing VAE (CDP-VAE), which uses windowing pooling filters. Evaluation using Fréchet Inception Distance (FID) and Mean Squared Error (MSE) shows that the Q-VAE outperforms both, suggesting its potential for image resolution enhancement and synthetic data generation.
@@ 1 Introduction
Variational Autoencoders (VAEs) have become foundational tools in generative modeling and image reconstruction, especially in tasks that require efficient and precise feature extraction from high-resolution data [1]. VAEs are part of the broader family of latent-variable models [2], wherein an encoder maps input data to a structured latent space, and a decoder reconstructs the data from this representation. Despite their success in a wide range of applications, VAEs often face challenges as latent space is often not expressive enough to model intricate relationships, often leading to blurry or low-quality output [3]. These limitations arise primarily from the difficulty of capturing complex high-resolution dependencies, which become especially apparent when dealing with large data sets (which consists of millions of images) or complex structures [5].
Recent advancements in quantum computing offer promising avenues for overcoming these challenges by exploiting quantum phenomena such as superposition and entanglement. These properties allow for more efficient data representation and processing, potentially overcoming some of the bottlenecks inherent in classical methods for generative modeling [6]. The integration of quantum circuits into neural networks, particularly within the VAE encoder, has shown potential to improve latent space representations, leading to enhanced feature extraction and more accurate reconstructions [7]. Quantum in VAE represents an example of such hybrid models, which leverage quantum operations to better capture the underlying data structure, making them especially well suited for high-resolution image tasks [6].
Recent advances in generative models have led to the increased use of low-resolution images, such as 8x8 or 16x16 pixels, as starting points for training medium-quality output [8]. This approach might seem counterintuitive because of the small size of these images, but it has proven effective for several reasons. Training generative models on low-resolution images allows for faster training times, as smaller images require less computational power and processing, enabling faster iterations and fine-tuning. In addition, low-resolution images force the model to focus on learning high-level patterns and structures rather than fine-grained details. This simplification helps the model build a strong foundation for understanding the essential features of the data. Once the model has learned to generate meaningful, broad structures at low resolution, it can then be used to upscale or refine the output to a medium resolution, such as 32x32, using techniques like super-resolution or progressive refinement. This method allows the model to generate more detailed and sharper images, while avoiding the computational difficulties that come with directly training on medium resolutions like 32x32.
Although 32x32-pixel images are commonly used in many applications, including image classification and generative modeling, classical models like VAEs struggle with these resolutions due to limitations in their architecture [9]. VAEs operate by encoding input data into a compressed probabilistic latent space that is then used to reconstruct the input image. However, with 32x32 images, the encoder has to condense a large amount of pixel information into a smaller latent space, often losing fine details in the process. This leads to blurred or artifact-filled reconstructions, as the latent space might not capture the complexity of the image at this resolution. The encoder’s limited capacity to represent detailed features in the latent space becomes particularly problematic for 32x32 images, which are still complex enough to contain important texture, color variation, and structural information. As a result, VAEs can struggle to generate high-quality images from these lower-resolution inputs [10]. The use of 32x32 resolution is prevalent in many machine learning applications, particularly in image classification tasks and generative modeling, as it strikes a balance between computational efficiency and sufficient detail. Datasets like CIFAR-10 and CIFAR-100, which consist of 32x32 pixel images, are standard benchmarks for evaluating generative models and classifiers [11]. Additionally, modern generative models often employ techniques like progressive training, where low-resolution images (8x8 or 16x16) are initially learned and then progressively refined to medium resolutions, such as 32x32 or beyond. This process stabilizes training and reduces the complexity associated with generating high-resolution outputs [12]. In practice, combining low-resolution training (e.g., 8x8 or 16x16) with upscaling techniques allows generative models to learn more efficiently and handle the complexities of higher-resolution images. By training on smaller images first, the models can focus on understanding the broader patterns in the data, then gradually increase the resolution to produce more detailed output. This method has potential to be useful for reducing computational load and improving the quality of the final images generated.
In this work, we introduce a novel Q-VAE that integrates quantum encoding techniques in the encoder while using Convolutional Neural Networks (CNNs) in the decoder. This hybrid architecture leverages the high expressiveness of quantum states to improve feature extraction and mitigate the blurriness often associated with classical VAEs. Unlike traditional quantum generative models, which typically operate on lower-resolution datasets (e.g. 8x8 or 16x16 pixels) [12], our approach is designed to handle medium-resolution inputs, such as 32x32 pixels, allowing the model to produce better-quality image reconstructions for more complex datasets. In the context of Q-VAE, we use MNIST and USPS datasets primarily due to their simplicity and relevance to early-stage exploration of quantum models in image generation. Both are grayscale datasets, which are less complex than color images, making them ideal for testing new techniques, particularly when working with emerging technologies like quantum computing in machine learning.
Since quantum computing, especially its integration into deep learning models, is still in the early stages, focusing on simpler datasets allows for a more controlled environment where we can better understand the impact of quantum techniques on model performance. Furthermore, the datasets used in this study, such as MNIST and USPS (with image resolutions of 28x28 and 16x16 pixels, respectively), are relatively small in size, which is advantageous when exploring the potential of quantum circuits for feature extraction and latent space representation. Their reduced resolution helps reduce the computational burden compared to higher-resolution images. To ensure a fair evaluation, the MNIST dataset is first down-sampled from 28x28 to 16x16 pixels, and then, like the USPS dataset, it is up-sampled to 32x32 pixels before being used as input to assess our model’s performance.
By focusing on grayscale images, we reduce the complexity of the problem, allowing for clearer insights into how quantum techniques can improve feature extraction and image reconstruction, particularly for simpler tasks like recognizing digits. Once the approach is successfully validated on these datasets, the methods can be extended to more complex and higher-resolution data, such as color images or higher resolutions.
Quantitative analysis has been performed to compare the effectiveness of the model using common performance metrics, such as the Fréchet Inception Distance (FID) and the Mean Squared Error (MSE), which are essential to assess the quality of generated images and the representation of the latent space [18, 19]. Thus, by leveraging these datasets, we can benchmark the Q-VAE against traditional models,to explore that if quantum component brings measurable improvements to the generative process. In summary the following contributions have been made in this paper:
@@
  • Introduces a new model quantum VAE (Q-VAE), where quantum encoding is applied only in the encoder, simplifying the model and improving the feature extraction capabilities.
@@
  • Compares Q-VAE with the classical Direct Passing CDP-VAE, highlighting performance differences due to the inclusion of quantum encoding in the encoder.
@@
  • Avoids quantum entanglements or layers in the encoder, focusing solely on an efficient quantum encoding circuit for latent space encoding.
@@
  • Minimizes quantum-trainable parameters, leading to more efficient model training and better utilization of quantum resources.
@@
  • Evaluate whether quantum encoding improves the generative quality and the latent space representation compared to classical methods.
@@
  • Investigates whether converting images from low resolution (16x16) to medium resolution (32x32) improves model performance.
@@
The structure of this paper is as follows. Section 2 provides a review of related work on classical VAEs and their integration with quantum computing. Section 3 presents the background on Variational Autoencoders (VAEs). In Section 4, the methodology for the proposed quantum-enhanced VAE model is introduced. Section 5 outlines the numerical simulations conducted, while Section 6 discusses the performance analysis. Finally, Section 7 concludes the paper, summarizing the key findings and suggesting potential avenues for future research.
@@ 2 Related Research
In recent years, variational auto-encoders (VAEs) have garnered significant attention in generative modeling, especially for image reconstruction tasks [21]. Traditional VAEs are praised for their ability to encode data into a structured latent space and regenerate approximations of the input. However, classical VAEs often suffer from blurred reconstructions, [3], prompting research to improve VAE architectures for better image quality and feature extraction [25, 26, 27]. Several studies have focused on enhancing VAE performance using convolutional architectures [43, 30, 31, 32]. Gatopoulos et al. introduced convolutional VAEs to better capture image features, surpassing the performance of fully connected architectures in medium-resolution data sets such as CIFAR-10 and ImageNet32 [28]. Furthermore, combining VAE with adversarial training, such as in Adversarial Autoencoders (AAE) [33], incorporates discriminator networks to enhance the visual quality of generated images, address the blurriness issues inherent in classical VAEs, and improve machine learning applications. In addition, anti-adversarial VAEs [34, 35] and VAE-GAN hybrids [36] enhance image generation by incorporating adversarial training, leading to sharper images by reducing blurriness [37].
Quantum Machine Learning (QML) has emerged as a promising field, with studies exploring the integration of quantum circuits into generative models. Quantum GANs (QGANs) leverage quantum computing to enhance the data representation and image generation capabilities of GANs. Although QGANs are not directly related to VAEs, they represent a quantum-based generative approach that improves data diversity and generation quality [39]. Quantum Variational Autoencoders [40, 41, 42, 43, 44, 45] build on the foundational work by Khoshaman et al., who proposed a quantum VAE where the latent generative process is modeled as a quantum Boltzmann machine (QBM). Their approach demonstrates that quantum VAE can be trained end-to-end using a quantum lower bound to the variational log-likelihood, achieving state-of-the-art performance on the MNIST dataset [38]. They have used Log-Likelihood (LL) and Evidence Lower Bound (ELBO) as the measurement and have shown quantum VAE confidence level for the numerical results is smaller than ±0.2 in all cases. Gircha et al. further advanced the quantum VAE architecture by incorporating quantum circuits in both the encoder, which utilizes multiple rotations, and the decoder [42]. These quantum VAE models primarily focus on leveraging quantum methods in the latent space representation, enhancing the generative capabilities of VAEs.
This paper addresses a gap in the literature where quantum techniques are primarily used in hybrid models for the encoder, while the decoder remains classical. Our proposed model is unique in that it leverages quantum encoding with a single rotation gate and measurement solely in the encoder, distinguishing it from other approaches.
@@ 3 Background
Variational Auto-encoders (VAEs) are widely recognized as powerful generative models in machine learning. They are based on a probabilistic framework where an encoder network maps input data $\mathbf{x}$ to a latent distribution $q(\mathbf{z}|\mathbf{x})$, and a decoder network reconstructs the data from the latent variables $\mathbf{z}$ as shown in equation 1 [4].
$L_{\theta,\phi}(x)=\mathbb{E}_{q_{\phi}(z|x)}\left[\log p_{\theta}(x,z)-\log q_{\phi}(z|x)\right]$ (1)
where $\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}$ represents the expectation with respect to the approximate posterior $q(\mathbf{z}|\mathbf{x})$. The reconstruction term encourages the model to generate data points that are similar to the observed data, while the KL divergence regularizes the learned latent variables by pushing the approximate posterior towards the prior distribution. Although VAEs have been successful in many domains, including image generation and anomaly detection, they often face significant challenges when applied to high-resolution data. This issue can arise because the latent space is often not expressive enough to model the intricate relationships. [14].
@@ 3.1 Encoder
A classical VAE encoder is typically implemented as a fully connected network or a Convolutional Neural Network (CNN), with the aim of mapping the input data to a lower-resolution latent space. The primary objective of the encoder is to process the input data $x$, transform it through a series of hidden layers, and output two parameters: the mean $\mu$ and the log variance $\log(\sigma^{2})$, which are used for variational inference [1]. In the case of image data, the input image is first reshaped into a vector and passed through multiple fully connected layers. The transformations performed by these layers help extract important features from the input. For example, given an input image of dimension 1024 (flattened from a $32\times 32$ image), the encoder progressively reduces the resolution’s, often mapping the input image $x_{0}$ into a lower-resolution space $x_{1}$.
3.1.1 CNN or Fully Connected Layers in Encoders
CNNs are particularly useful for image-based data because they can capture local spatial patterns (e.g., edges, textures) by using convolution layers. These convolution layers progressively down-sample the input image and extract high-level features, while the fully connected layers act to compress these features into a more compact latent representation. In CNN-based encoders, after the convolution layers extract relevant patterns, fully connected layers are used to map the high-resolution feature maps into the latent space. This approach allows the model to better capture spatial dependencies while still benefiting from the representational power of fully connected layers.